# State-adaptive quantum error correction and fault-tolerant quantum computing

### v1

@@ Abstract
We present a theoretical framework for state-adaptive quantum error correction (SAQEC) that bridges the gap between quantum computing and error correction paradigms. By incorporating knowledge of quantum states into the error correction process, we establish a new capacity regime governed by quantum mutual information rather than coherent information. This approach reveals a fundamental connection to entanglement-assisted protocols. We demonstrate practical applications in fault-tolerant quantum computation, showing how state-adaptivity enables enhanced error correction without additional measurement overhead. The framework provides new insights into quantum channel capacities while offering implementation advantages for current quantum computing platforms.
@@ I Introduction
An unknown classical bit string can be easily determined and cloned. For a quantum state, being known or unknown makes a huge difference, highlighted by the no-cloning theorem WZ82; Die82. The standard quantum error correction (QEC) theory requires the decoupling of environmental degree of freedom (e.g. eavesdropper) from the encoded information, hence forbidding cloning and making quantum capacity being private KL97; Dev05; HHW+08; Kle07. However, we realize that the decoupling and privacy is unnecessary, and it is possible to establish a new setting for QEC by making the encoded source information visible, i.e., a state-adaptive setting. We find that, as expected, this is the quantum analog of the classical channel capacity measured by mutual information.
In this work, we establish the state-adaptive (SA) QEC theory. We show that the SA quantum capacity is provided by the quantum mutual information by both an equivalence with the entanglement-assisted (EA) model and also a direct proof. The equivalence highlights the role of entanglement and quantum teleportation BBC+93: an unknown state can be teleported consuming entanglement, while a known state can be directly prepared according to its classical description. On the dual side, the quantum dense coding BW92 can also be substituted by SA schemes for sending classical information over quantum channels. Therefore, we obtain the capacity diagram in Fig. 1. By comparison, the EA model is more suitable for quantum communication whereas the assisted entanglement can be pre-established, while the SA model is more suitable for quantum computing whereas the knowledge of the states can benefit error correction.
@@
Figure 1: The diagram to show the relation among the coding models and their capacities, with $C_{H}\geq P\geq Q$, $C_{EA}\geq C_{H}$, $Q_{EA}\geq Q$, and $C_{EA}=2Q_{EA}$. Here $C_{H}$, $P$, $Q$ is the Holevo, private, and quantum capacity of a quantum channel, respectively.
@@ V Direct proof of state-adaptive quantum capacity
To gain a better understanding of the state-adaptive setting, we provide a direct proof of the state-adaptive quantum capacity. Following standard approaches Hay17; Wil17; Wat18, the method for the decoding is also using a Petz map Pet86; OP93; Wil17; BK02; NM10; GLM+22. Given a state $\rho$ and a channel $\Phi$, the Petz map $\mathcal{R}_{\rho,\Phi}$ is specified by the set of Kraus operators $\{R_{i}\}$ with
@@ VI Relation among capacities
Classical SA quantum Holevo Private Quantum
Measure $I(X:Y)$ $I(\rho_{A}:\rho_{B})/2$ $I(X:\rho_{B})$ I(X:Y)I(X:E)I(X:Y)-I(X:E) $I(\rho_{A}:\rho_{B})-I(\rho_{A}:\rho_{E})$
Additivity sub sub super super super
Decoupling no no no yes yes
Table 1: Channel capacities with key features. It shows the (SA) quantum capacity of a quantum channel as the analog of (classical) private capacity of a classical channel, while the Holevo setting is a hybrid model, which does not require a fully decoupling of the environment DST14.
The state-adaptive model fits into the rich family of quantum channel coding models. This is shown in Table 1. The SA quantum capacity is the analog of classical capacity, both of which are not private, and the usual quantum capacity is the analog of private capacity, and the one in the middle is the Holevo model.
The knowledge of the input data source $\rho$ is powerful. It is easy to see there is no need to use EA in the state-adaptive model since given the knowledge of the input source $\rho$, encoding $V$, and also the noise channel $\Phi$, the Petz map gives the optimal recovery BK02; NM10. For the proof, only a partial Petz recovery is needed to correct typical errors. The role of EA is understood from quantum teleportation and its extension: an unknown state $\rho$ can be transmitted by sending the classical bits from Bell measurements with assisted Bell states, while optimal assisted state can be constructed which factorizes into a sum of Bell states each in a typical subspace Wil17. Feedback would not increase the capacity $Q_{SA}$ neither for the same reason, and indeed it has been proven that feedback cannot change $C_{EA}$ Bow04.
@@ VII Transversal fault-tolerant universal quantum computing via coding switching
The major class of quantum codes are stabilizer codes Got98. A stabilizer code is a code specified by a stabilizer group. Namely, a $[[n,k,d]]$ stabilizer code $\mathcal{C}$ encoding $k$ logical qubits into $n$ physical qubits is defined by nkn-k independent commuting stabilizers $\langle s_{i}\rangle$, which generates an abelian subgroup of the $n$-qubit Pauli group. The decoder is to measure its commuting stabilizers by projectors $P_{i}=(\mathds{1}+s_{i})/2$, and the identified errors are effectively Pauli errors, no matter what the physical noise is since a Kraus operator can be expanded as a superposition of Pauli errors. In this sense, it is neither channel-adaptive nor state-adaptive. A gate $U$ is logical if it commutes with the code projector $P=\prod_{i}P_{i}$ for $[U,P]=0$, and for a stabilizer code, the elementary logical gates $X_{L}$ and $Z_{L}$ are usually transversal product of Pauli operators.
@@
For physical implementation, the scheme above can be adopted to various platforms. There are notable differences between solid-state platforms and the photonic platform. It is relatively easier for solid-state platforms, but for photonic platform with qubits encoded in discrete variable, the CNOT gate is probabilistic PJF01; BR05. A scalable approach to resolve this is the fusion-based scheme BBB21, which employs fusion measurements to grow a graph state efficiently. The graph state carries the information of the circuit to be executed, the encoding and code switching. One direction through the graph is the simulated time of the circuit, and the part of the graph at a later time can be generated gradually. The correlators of the graph can be used for error correction, which can correct more errors than using stabilizers of a code.
@@
VIII Conclusion
In this work, we proposed a framework of state-adaptive quantum error correction (SAQEC) which is a quantum analog of classical error correction. We showed that standard stabilizer codes can be employed in the setting of SAQEC, which can achieve higher coding rate and error threshold, without a significant overhead. By definition, an error threshold is the value of $\lambda$ at which a capacity of a parameterized channel $\Phi(\lambda)$ becomes zero. It is clear to see SAQEC allows a higher error threshold than the standard QEC.
In all, our finding is rather surprising. It resolves a longstanding question in the field: why quantum mutual information serves as the capacity measure for the entanglement-assisted coding model, which does not have classical analog. Our result emphasizes the importance of pre-knowledge of quantum states and operations, and may also benefit the study of the superadditivity phenomena in Shannon theory.
IX Acknowledgement
This work has been funded by the National Natural Science Foundation of China under Grants 12447101 and 12105343.

### v2

@@ Abstract
We present a theoretical framework for state-adaptive quantum error correction that bridges the gap between quantum computing and error correction paradigms. By incorporating knowledge of quantum states into the error correction process, we establish a new capacity regime governed by quantum mutual information rather than coherent information. This approach reveals a fundamental connection to entanglement-assisted protocols. We demonstrate practical applications in fault-tolerant quantum computation, showing how state-adaptivity enables enhanced error correction without additional measurement overhead. The framework provides insights into quantum channel capacities while offering implementation advantages for current quantum computing platforms.
@@ I Introduction
An unknown classical bit string can be easily determined and cloned. For a quantum state, being known or unknown makes a huge difference, highlighted by the no-cloning theorem WZ82; Die82. The standard quantum error correction (QEC) theory requires the decoupling of environmental degree of freedom (e.g. eavesdropper) from the encoded information, hence forbidding cloning and making quantum capacity be private KL97; Dev05; HHW+08; Kle07. However, we realize that the decoupling and privacy are unnecessary, and it is possible to establish a new setting for QEC by making the encoded source information visible, i.e., a state-adaptive setting. We find that, as expected, this is the quantum analog of the classical channel capacity measured by mutual information.
In this work, we establish the state-adaptive (SA) QEC theory. We show that the SA quantum capacity is provided by the quantum mutual information, as evidenced by its equivalence to the entanglement-assisted (EA) model and a direct proof. The equivalence highlights the role of entanglement and quantum teleportation BBC+93: an unknown state can be teleported consuming entanglement, while a known state can be directly prepared according to its classical description. On the dual side, the quantum dense coding BW92 can also be substituted by SA schemes for sending classical information over quantum channels. Therefore, we obtain the capacity diagram in Fig. 1. By comparison, the EA model is more suitable for quantum communication scenarios where assisted entanglement can be pre-established, whereas the SA model is better suited for quantum computing contexts where prior knowledge of quantum states facilitates error correction.
@@
Figure 1: The diagram to show the relation among the coding models and their capacities, with $C_{H}\geq P\geq Q$, $C_{EA}\geq C_{H}$, $Q_{EA}\geq Q$, and $C_{EA}=2Q_{EA}$. Here $C_{H}$, $P$, $Q$ are the Holevo, private, and quantum capacity of a quantum channel, respectively.
@@ V Direct proof of state-adaptive quantum capacity
To further elucidate the state-adaptive setting, we provide a direct proof of the state-adaptive quantum capacity, shown in the appendix which also contains examples for a few channels. Following standard approaches Hay17; Wil17; Wat18, the method for the decoding also uses a Petz map Pet86; OP93; Wil17; BK02; NM10; GLM+22. Given a state $\rho$ and a channel $\Phi$, the Petz map $\mathcal{R}_{\rho,\Phi}$ is specified by the set of Kraus operators $\{R_{i}\}$ with
@@ VI Relation among capacities
The knowledge of the input data source $\rho$ is powerful. It is easy to see there is no need to use EA in the state-adaptive model since given the knowledge of the input source $\rho$, encoding $V$, and also the noise channel $\Phi$, the Petz map gives the optimal recovery BK02; NM10. For the proof, only a partial Petz recovery is needed to correct typical errors. The role of EA is understood from quantum teleportation and its extension: an unknown state $\rho$ can be transmitted by sending the classical bits from Bell measurements with assisted Bell states, while optimal assisted state can be constructed which factorizes into a sum of Bell states each in a typical subspace Wil17. Feedback would not increase the capacity $Q_{SA}$ either for the same reason, and indeed it has been proven that feedback cannot change $C_{EA}$ Bow04.
@@ VII Transversal fault-tolerant universal quantum computing via coding switching
The major class of quantum codes are stabilizer codes Got98. A stabilizer code is a code specified by a stabilizer group. Namely, a $[[n,k,d]]$ stabilizer code $\mathcal{C}$ encoding $k$ logical qubits into $n$ physical qubits is defined by nkn-k independent commuting stabilizers $\langle s_{i}\rangle$, which generate an Abelian subgroup of the $n$-qubit Pauli group. The decoder is to measure its commuting stabilizers by projectors $P_{i}=(\mathds{1}+s_{i})/2$, and the identified errors are effectively Pauli errors, no matter what the physical noise is since a Kraus operator can be expanded as a superposition of Pauli errors. In this sense, it is neither channel-adaptive nor state-adaptive. A gate $U$ is logical if it commutes with the code projector $P=\prod_{i}P_{i}$ for $[U,P]=0$, and for a stabilizer code, the elementary logical gates $X_{L}$ and $Z_{L}$ are usually transversal product of Pauli operators.
@@
VIII Conclusion
In all, we proposed a framework of state-adaptive quantum error correction (SAQEC) which is a quantum analog of classical error correction. We showed that standard stabilizer codes can be employed in the setting of SAQEC, which can achieve higher coding rate and error threshold, without a significant overhead. By definition, an error threshold is the value of $\lambda$ at which a capacity of a parametrized channel $\Phi(\lambda)$ becomes zero. It is clear to see SAQEC allows an error threshold higher than that of the standard QEC.
Our finding is rather surprising. It resolves a longstanding question in the field: why quantum mutual information serves as the capacity measure for the entanglement-assisted coding model, which does not have the classical analog. Our result emphasizes the importance of pre-knowledge of quantum states and operations, and may also benefit the study of the superadditivity phenomena in Shannon theory.
IX Acknowledgement
This work has been funded by the National Natural Science Foundation of China under Grants 12447101 and 12105343. Many thanks to R. Laflamme for his insights on quantum error correction.
Appendix
.1 Channel capacity
The state-adaptive (SA) model fits into the rich family of quantum channel coding models. This is shown in Table 1. The state-adaptive quantum capacity is the analog of classical capacity, both of which are not private, and the usual quantum capacity is the analog of private capacity, and the one in the middle is the Holevo model.
In the classical setting the encoding is from bits to bits, with a source entropy $H(X)$, similarly for the state-adaptive setting which has a source entropy $S(\rho)$. In the Holevo setting the encoding is from bits to qubits, still with a source entropy $H(X)$, but each bit string $x$ is mapped to a state $\rho_{x}$ with $x$ as its label, and measurement is required to recover $x$ from $\rho_{x}$. The Holevo capacity is the mutual information of a classical-quantum state, which is not subadditive in general, while does not require a fully decoupling of the environment DST14. The quantum capacity, and also the private capacity of either a classical or a quantum channel, all require the decoupling of the environment state.
Measure Additivity Decoupling
Classical $I(X:Y)$ Sub No
SA quantum $I(\rho_{A}:\rho_{B})/2$ Sub No
Holevo $I(X:\rho_{B})$ Super No
Private I(X:Y)I(X:E)I(X:Y)-I(X:E) Super Yes
Quantum $I(\rho_{A}:\rho_{B})-I(\rho_{A}:\rho_{E})$ Super Yes
Table 1: Channel capacities with key features. It shows the (SA) quantum capacity of a quantum channel as the analog of (classical) private capacity of a classical channel, while the Holevo setting is a hybrid model.
Below is a direct proof of the state-adaptive quantum capacity.
Proof.
.2 Examples of noise channels
Besides Pauli channels which are common in the study of quantum error correction, there are also a few other examples that are practically relevant Wil17. We focus on the qubit case. The erasure channel $\mathcal{E}$ erases a qubit with error rate $p$ replacing it with a state $\rho_{e}$ that is outside the state space
$\mathcal{E}(\rho)=(1-p)\rho+p\rho_{e}.$ (9)
A notable fact in QEC is that correcting $t$ errors is equivalent to correcting $2t$ erasures as their locations can be identified. Its quantum capacity is $Q(\mathcal{E})=1-2p$, and SA quantum capacity is $Q_{SA}(\mathcal{E})=1-p$.
The amplitude-damping (AD) channel $\mathcal{A}$ is specified by two Kraus operators $A_{0}=|0\rangle\langle 0|+\sqrt{1-\gamma}|1\rangle\langle 1|$, $A_{1}=\sqrt{\gamma}|0\rangle\langle 1|$. It describes the decay from an excited state $|1\rangle$ to the ground state $|0\rangle$, and can be viewed as a ‘partial’ erasure channel. Actually, when encoded into two qubits as degenerate states $|0\rangle\equiv|01\rangle$, $|1\rangle\equiv|10\rangle$, the AD channels on the two qubits will act as an erasure channel with error rate $\gamma$. Code design for the AD channel can then be converted to that for erasure. Its quantum capacity is $Q(\mathcal{A})=\max_{a}h(a(1-\gamma))-h(a\gamma)$, and SA quantum capacity is $Q_{SA}(\mathcal{A})=\frac{1}{2}\max_{a}h(a)+h(a(1-\gamma))-h(a\gamma)$, for $a\in[0,1]$, $h(x)$ as the binary entropy function. The threshold value for $Q(\mathcal{A})$ is $\gamma=\frac{1}{2}$, while $Q_{SA}(\mathcal{A})$ is already nonnegative for any $\gamma$.
The qubit Pauli channel $\mathcal{P}$ is a random mixture of Pauli operators $\sigma_{i}=\{X,Y,Z\}$ and identity $\mathds{1}$ with probabilities $p_{x}$, $p_{y}$, $p_{z}$, and $p_{0}=1-p_{x}-p_{y}-p_{z}$. When $p_{x}=p_{y}=p_{z}$, it reduces to the depolarizing channel $\mathcal{D}$, and when $p_{x}=p_{y}=0$, it reduces to the dephasing channel. Among them, the quantum capacities for $\mathcal{P}$ ad $\mathcal{D}$ do not have a closed form due to the superadditivity of coherent information. Here we discuss the realization of SA capacity.
Finding capacity-approaching codes is highly nontrivial even when a channel capacity is computable. Here we study code design for the erasure channel due to its simplicity and its ‘universal’ property, namely, other channels can be converted to erasure channels with a suitable encoding. With the three-qubit repetition code $|0\rangle\equiv|000\rangle$, $|1\rangle\equiv|111\rangle$, measuring the two stabilizers $Z_{1}Z_{2}$, $Z_{2}Z_{3}$ can define an erasure event: any nontrivial syndrome is treated as an effective erasure. For $\mathcal{D}$ with an error rate $p$, the effective erasure channel has an error rate $1-(1-p)^{3}\geq p$. This is the analog of the two-qubit encoding for the AD channel above.
For an erasure channel $\mathcal{E}$ with error rate $p$, the quantum capacity $Q(\mathcal{E})=1-2p$ as otherwise it would violate the no-cloning theorem BDS97. In the EA model, it is increased to $Q_{EA}(\mathcal{E})=1-p$, which means even when $p>\frac{1}{2}$, the assisted entanglement will help for the transmission of qubits. For the SA model, we shall use the knowledge of the input and the channel. Note that $\mathcal{E}$ with $p$, denoted by $\mathcal{E}_{p}$, can be simulated by
$\mathcal{E}_{p}=\frac{1}{2}(\mathcal{E}_{2p-1}+\mathcal{E}_{1})$ (10)
for $p>\frac{1}{2}$. Let $p_{1}=2p-1$, then $\mathcal{E}_{p_{1}}$ can be simulated again via $\mathcal{E}_{p_{2}}$ for $p_{2}=2p_{1}-1$ and $\mathcal{E}_{1}$. This concatenation carries on and yields $p_{r}=1-2^{r}(1-p)$. Requiring $p_{r}\in[0,\frac{1}{2}]$, this implies $p\in[1-2^{-r},1-2^{-r-1}]$ approaches the maximal value 1.
This leads to an SA coding scheme: with $\mathcal{E}_{p}$ ($p\in[\frac{1}{2},1]$) simulated by the concatenation at level $r$, a code designed for $\mathcal{E}_{p_{r}}$ ($p_{r}\in[0,\frac{1}{2}]$) is used as an outer code, and a $r$-level concatenated repetition code is used as an inner code, so that the source is input to $n$ parallel uses of the channels $\mathcal{E}_{p_{r}}$, while other channels $\mathcal{E}_{1}$ only output junk data. It is easy to see the coding rate approaches 1p1-p, exponentially fast with $r$. In all, this forms an SA coding scheme simulated by the usual quantum coding scheme for $\mathcal{E}_{p}$ ($p\in[0,\frac{1}{2}]$), with a structured input source and channel, treating the ‘bad’ channels $\mathcal{E}_{1}$ as free resources. Note that the division of channels into ‘good’ and ‘bad’ ones echoes the method of polar codes RL09.
There are very noisy channels that cannot be converted into erasure channels, though. The class of entanglement-breaking (EB) channels are such examples HSR03. An EB channel $\mathcal{B}$ is specified by a positive operator-valued measure (POVM) $\{M_{i}\}$ and a set of states $\{\rho_{i}\}$ as
$\mathcal{B}(\rho)=\sum_{i}\text{tr}(M_{i}\rho)\rho_{i},$ (11)
and its quantum capacity is zero, $Q(\mathcal{B})=0$, while $Q_{EA}(\mathcal{B})=C_{H}(\mathcal{B})=\chi(\mathcal{B})$, the optimal Holevo mutual information Wil17. It is a one-shot form since entanglement does not help for the coding. For the SA model, a source $k$-qubit state $\rho$ could be entangled, which will be destroyed by EB channels. This implies that one has to send the classical description $\rho]$ over the EB channels, which then induces $Q_{SA}(\mathcal{B})$ being $\chi(\mathcal{B})$. There exists no SA coding that can directly protect quantum state $\rho$ over EB channels. In other words, when $Q$ is exactly zero for a channel, it cannot be boosted to achieve $Q_{SA}$ in the SA setting. On the contrary, in the EA setting which is more suitable for communication, the boosted capacity is due to the assisting entanglement that needs to be pre-established and maintained noiseless.
.3 Physical implementation
@@ IX Acknowledgement
Here we discuss platforms for physical implementation. There are notable differences between solid-state platforms and the photonic platform. It is relatively easier for solid-state platforms, but for photonic platform with qubits encoded in discrete variable, the CNOT gate is probabilistic PJF01; BR05. A scalable approach to resolve this is the fusion-based scheme BBB21, which employs fusion measurements to grow a graph state efficiently. The graph state carries the information of the circuit to be executed, the encoding and code switching. One direction through the graph is the simulated time of the circuit, and the part of the graph at a later time can be generated gradually. The correlators of the graph can be used for error correction, which can correct more errors than using stabilizers of a code.